OpenStudy (anonymous):

sketch the graph y =1/x+4

OpenStudy (anonymous):

\[\frac{1}{x+4} \ or \ \frac{1}{x} + 4\]

OpenStudy (anonymous):

how does the graph go

OpenStudy (anonymous):

They are both hyperbolas, with different asymptotes and centers

OpenStudy (anonymous):

that i know, but what governs the direction of these hyberbolas and from where do they start

OpenStudy (anonymous):

What I did was to find a graphing tool to make a plot. What the "sketch" part of the question means is to have an idea about the form from the equation, and find a few points that fit the curve. You still haven't answered my very first question?

OpenStudy (anonymous):

Assuming it's the second one \[y = \frac{1}{x} + 4\] \[xy = 1 + 4x\] So these are not "east-west" or "north-south" but angled like xy=1 Notice that x can never be equal to zero y can never be equal to 4